# Disclosing the catalog pulsars dominating the Galactic positron flux

### v1

@@ 2.1 ATNF catalog
The sample of pulsars considered in this work is taken from the ATNF pulsar catalog [42], version 2.1.1  [^1] http://www.atnf.csiro.au/research/pulsar/psrcat, which lists the observed rotation-powered pulsars, including those detected only at high energies. Within this version of the ATNF catalog, the distance to each pulsar is estimated with the most updated electron-density model [43]. Sources are selected by excluding all the millisecond pulsars, namely cutting off rotation periods $P_{0}<0.01$ s. The emission coming from these sources would require a different model with respect to what implemented in this work. Furthermore, it has been shown to be negligible through an analysis realized with simulations of the millisecond pulsar population [44].
@@ 2.2 Positron emission model
through which we obtain the value of $L_{0}$, fixing $E_{1}$=0.1 GeV [48, 49]. The parameter $\eta$ encodes the efficiency of conversion of the spin-down energy into $e^{+}$ (which is half of the efficiency of conversion into $e^{\pm}$). $W_{0}$ is the initial rotational energy of a pulsar and can be computed from the pulsar age $T$, the spin-down luminosity $\dot{E}$ and the decay time $\tau_{0}$:
@@ 2.2.1 Simulated emission parameters
where $n_{0}$ is the ISM density around the SNR taken to be 3 cm-3, $E_{SN}$ = $10^{51}$ erg is the energy emitted by the SN explosion and $v_{k}$ is the birth velocity of the pulsar. The formalism is reported in [56], to which we refer for further details. To test the scenario described by Eq. (2.6), for each source its birth-kick velocity is simulated, adopting its distribution as reported in [57] (FK06VB) and implemented in gammapy.astro.population [52, 53], which is the sum of two Gaussians (see their Eq. 7) for each of the three velocity components.
@@ 2.3 Positron propagation to the Earth and anisotropy
The propagation parameters are fixed to the ones derived in [6] from a fit to the latest data measured by the AMS-02 experiment, particularly the absolute fluxes of protons, He, C, O, N, B/C, Be/C, and Li/C after 7 years of data collection, from 2011 to 2018 [60], complemented with proton and helium data from Voyager [61]. Specifically, we consider their benchmark model Conv $dv_{c}/dz$, with the value of $L$ fixed at $4$ kpc (see Table 2 of [6] for the values of the parameters of this model). We note that, however, the parameter $L$ is not relevant in this study since we implement solutions without boundaries both in the radial and the vertical directions. The infinite halo approximation has been widely used to compute the flux from single sources located in the Galactic plane in [36, 62, 40].
@@ 3.1 Results from fits to AMS-02 $e^{+}$ data
In the Delayed injection setup, only the $e^{+}$ produced after the pulsar has left the remnant are considered. This approach reduces the influence of extremely bright sources, as most $e^{+}$ are produced in the early stages of the pulsar’s lifetime, i.e. $\lesssim 50$ kyr. Consequently, the total flux has not strong peaks and follows generally a smooth trend. However, the fit requires efficiencies between 5 and 46% - implied in the value of $A_{P}$ since the model does not account for the significant amount of $e^{+}$ produced at the beginning of the pulsar’s lifetime. Efficiencies close to 50%, although compatible with previous estimates obtained with the same model [41, 39], are in tension with the values obtained from the analysis of $\gamma$-ray halos around pulsars [16, 30, 77, 32], assuming a slow diffusion around this sources. They result although to be compatible with the ballistic scenario proposed in [30], which does not invoke a suppressed diffusion around pulsars. In all the three models, good fits come with a normalization $A_{S}$ of secondary $e^{+}$ close to one, which marks a significant improvement compared to previous predictions [41, 4]. We ascribe this result to the new computation of the secondary source term, in which the $e^{+}$ production cross sections have been adapted to a bunch of accelerator data [7].
@@ 3.3 The anisotropy of the $e^{+}$ flux
The anisotropy of the $e^{+}$ flux could in principle be a valid complementary observable in order to understand some physical properties of the emitting sources [35]. We show here the results of the total $e^{+}$ dipole anisotropy resulting from the contribution of all sources for the realizations that produce a good fit to the data for $\mathbf{\tau_{0}}$ distribution, $\mathbf{\tau_{0}}$ fixed and Delayed injection models. We compute the anisotropy for each realization, and provide the average value together with the 68% containment band. We predict an anisotropy level between $10^{-4}$ and $10^{-3}$ in the whole energy range. $\mathbf{\tau_{0}}$ distribution, which is the model that more frequently present the domination of a few pulsars, produce a higher anistropy level with respect to the others, but still well below the AMS-02 upper limits of $\Delta_{e^{+}}$ obtained from the 10 years data [88], which stand at least two orders of magnitude above our predictions. It is therefore evident that the properties of ATNF pulsars as emerging from AMS-02 flux data are very unlikely be tested by present or forthcoming data on the $e^{+}$ anisotropy. Our results are similar to the ones found in [35], but with an updated comparison to available experimental upper limits, and using model tuned to the recent insights on $e^{\pm}$ emission coming from multi-wavelength observations.
@@ 4.1 The "three musketeers"
According to the latest version of the ATNF catalog, the pulsar B1055-52 (J1057-5226) is located at a distance of 0.093 kpc. However, the true distance to B1055-52 remains uncertain, varying between 714 and 93 pc depending on the Galactic free-electron density models used: [90] and [91], respectively, despite both using the same dispersion measure. This pulsar is highly energetic, with a spin-down power of $3\times 10^{34}$ erg/s and a spin-down age of 535 kyr. Alongside Geminga and Monogem, B1055-52 is one of the most studied nearby middle-aged pulsars, showing $\gamma$-ray, X-ray, and radio emissions, as highlighted in a recent multi-wavelength study [92]. The pulsar is associated with the point-like $\gamma$-ray source 4FGL J1057.9-5227 in the Fermi-LAT 4FGL catalog [93], showing emissions up to a few GeV. No TeV emissions have been detected thus far. A significant X-ray flux has been observed, including in the recent eRASS:4 catalog obtained by SRG/eROSITA in the 0.2-2.3 keV energy band [94]. Although prominent PWN emission has not been detected in the X-ray band, faint extended emission in a small annulus around the pulsar, possibly linked to its PWN, has been observed by Chandra [95]. The possible detection of a $\gamma$-ray halo around this pulsar is very challenging with GeV data from Fermi-LAT because, due to the proximity of the source, the halo could span several tens of degrees or even cover the entire sky. Detecting a signal in X-ray data is also likely prohibitive due to the typically small field of view of detectors. However, at multi-TeV energies, this source can be detected as an extended source a few degrees in size. It is not within the field of view of HAWC or LHAASO, but it can be detected by HESS.

### v2

@@ 2.1 ATNF catalog
The sample of pulsars considered in this work is taken from the ATNF pulsar catalog [42], version 2.1.1  [^1] http://www.atnf.csiro.au/research/pulsar/psrcat, which lists the observed rotation-powered pulsars, including those detected only at high energies. Within this version of the ATNF catalog, the distance to each pulsar is estimated using the most updated electron-density model [43]. However, the choice of electron-density model introduces unavoidably biases in the distance estimates, which can significantly impact the computation of the $e^{+}$ flux. The model in [43] could lead to artificial small distances of some pulsars, as for instance the distance to the source B1055-52, which can vary between 714 pc and 93 pc, depending on the Galactic free-electron density model used [44, 43].
Sources are selected by excluding all the millisecond pulsars, namely cutting off rotation periods $P_{0}<0.01$ s. The emission coming from these sources would require a different model with respect to what implemented in this work. Furthermore, it has been shown to be negligible through an analysis realized with simulations of the millisecond pulsar population [45].
@@ 2.2 Positron emission model
through which we obtain the value of $L_{0}$, fixing $E_{1}$=0.1 GeV [49, 50]. The parameter $\eta$ encodes the efficiency of conversion of the spin-down energy into $e^{+}$ (which is half of the efficiency of conversion into $e^{\pm}$). Changing the value of $E_{1}$ affects the normalization of the total spectrum, which in turn impacts the efficiency $\eta$ of individual sources. For the injection spectrum parameters used in this work, as listed in Table 1, setting $E_{1}$ to the $e^{+}$ mass can result in a maximum renormalization of the flux obtained with $E_{1}=0.1$ GeV by a factor of 0.65. This occurs for $\gamma_{L}=2$ and $\gamma_{F}=2.8$, which represents the specific parameter set that maximizes the effect of lowering the $E_{1}$ value. Choosing, on the other hand, values of $E_{1}$ larger than 0.1 GeV would lead to a flux renormalization greater than 1, thereby requiring lower efficiencies. $W_{0}$ is the initial rotational energy of a pulsar and can be computed from the pulsar age $T$, the spin-down luminosity $\dot{E}$ and the decay time $\tau_{0}$:
@@ 2.2.1 Simulated emission parameters
where $n_{0}$ is the ISM density around the SNR taken to be 3 cm-3, $E_{SN}$ = $10^{51}$ erg is the energy emitted by the SN explosion and $v_{k}$ is the birth velocity of the pulsar. The formalism is reported in [57], to which we refer for further details. The parameter values are taken from [39] to ensure consistency with the analysis framework used in that study. To test the scenario described by Eq. (2.6), for each source its birth-kick velocity is simulated, adopting its distribution as reported in [58] (FK06VB) and implemented in gammapy.astro.population [53, 54], which is the sum of two Gaussians (see their Eq. 7) for each of the three velocity components.
@@ 2.3 Positron propagation to the Earth and anisotropy
The propagation parameters are fixed to the ones derived in [6] from a fit to the latest data measured by the AMS-02 experiment, particularly the absolute fluxes of protons, He, C, O, N, B/C, Be/C, and Li/C after 7 years of data collection, from 2011 to 2018 [61], complemented with proton and helium data from Voyager [62]. Specifically, we consider their benchmark model Conv $dv_{c}/dz$, with the value of $L$ fixed at $4$ kpc (see Table 2 of [6] for the values of the parameters of this model). We note that, however, the parameter $L$ is not relevant in this study since we implement solutions without boundaries both in the radial and the vertical directions reported in Eq. 2.7. The infinite halo approximation has been widely used to compute the flux from single sources located in the Galactic plane in [36, 63, 40]. We verified that our conclusions remain unchanged when adopting a more refined approach, incorporating the halo size $L=4$ kpc using the image charge method described in [64], which modifies the infinite halo solution reported in our Eq. 2.7 (see Eq. 2 of their paper). By evaluating 100 realizations of one of our setups, later referred to as the $\bf\tau_{0}$ distribution model (see Sec. 2.4), we found that the mean deviation between the total $e^{+}$ flux from pulsars obtained with the boundary condition and the infinite halo approximation is less than 1% at $E=10$ GeV and decreases further at higher energies.
@@
The analysis is based on the standard, homogeneous diffusion model of CR propagation, without accounting for additional uncertainties that may arise from more refined propagation models. This standard model is calibrated to reproduce observations of diffuse CR nuclei fluxes but does not include localized effects near Earth. Such effects could be especially relevant for $e^{+}$ and $e^{-}$ fluxes, given their relatively shorter diffusion-loss length with respect to stable CR nuclei. In particular, this approach overlooks the influence of nearby structures like the Local Bubble [82], superbubbles, and the orientation of the regular Galactic magnetic field, which can induce anisotropic diffusion. Models of the Galactic magnetic field, such as those described in [83, 84, 85], suggest that anisotropic diffusion can significantly alter the contribution of individual sources to the observed flux at Earth. While these considerations are beyond the scope of the current study, they underscore the interest to account for such effects, characterized by large uncertainties, in future analyses to achieve a more complete understanding of CR propagation and $e^{\pm}$ fluxes.
@@ 3.1 Results from fits to AMS-02 $e^{+}$ data
In the Delayed injection setup, only the $e^{+}$ produced after the pulsar has left the remnant are considered. This approach reduces the influence of extremely bright sources, as most $e^{+}$ are produced in the early stages of the pulsar’s lifetime, i.e. $\lesssim 50$ kyr. Consequently, the total flux has not strong peaks and follows generally a smooth trend. However, the fit requires efficiencies between 5 and 46% - implied in the value of $A_{P}$ since the model does not account for the significant amount of $e^{+}$ produced at the beginning of the pulsar’s lifetime. Efficiencies close to 50%, although compatible with previous estimates obtained with the same model [41, 39], are in tension with the values obtained from the analysis of $\gamma$-ray halos around pulsars [16, 30, 79, 32], assuming a slow diffusion around this sources. They result although to be compatible with the ballistic scenario proposed in [30], which does not invoke a suppressed diffusion around pulsars. Changing the value of $E_{1}$ in the normalization of the spectrum from 0.1 to the $e^{+}$ mass can, for some specific set of parameters in the injection spectrum ($\gamma_{L}$ close to 2 and $\eta>33$% of the single pulsars), lead to unphysical values of $\eta$ for individual sources. However, this occurs only in a small region of the parameter space. The average value of $\eta$ required remains below 50%. In all the three models, good fits come with a normalization $A_{S}$ of secondary $e^{+}$ close to one, which marks a significant improvement compared to previous predictions [41, 4]. We ascribe this result to the new computation of the secondary source term, in which the $e^{+}$ production cross sections have been adapted to a bunch of accelerator data [7].
@@ 3.3 The anisotropy of the $e^{+}$ flux
The anisotropy of the $e^{+}$ flux could in principle be a valid complementary observable in order to understand some physical properties of the emitting sources [35]. We show in Fig. 8 the results of the total $e^{+}$ dipole anisotropy resulting from the contribution of all sources for the realizations that produce a good fit to the data for $\mathbf{\tau_{0}}$ distribution, $\mathbf{\tau_{0}}$ fixed and Delayed injection models. We compute the anisotropy for each realization, and provide the average value together with the 68% containment band. We predict an anisotropy level between $10^{-4}$ and $10^{-3}$ in the whole energy range. $\mathbf{\tau_{0}}$ distribution, which is the model that more frequently present the domination of a few pulsars, produce a higher anistropy level with respect to the others, but still well below the AMS-02 upper limits of $\Delta_{e^{+}}$ obtained from the 10 years data [94], which stand at least two orders of magnitude above our predictions. It is therefore evident that the properties of ATNF pulsars as emerging from AMS-02 flux data are very unlikely be tested by present or forthcoming data on the $e^{+}$ anisotropy. Our results are similar to the ones found in [35], but with an updated comparison to available experimental upper limits, and using model tuned to the recent insights on $e^{\pm}$ emission coming from multi-wavelength observations.
@@ 4.1 The "three musketeers"
According to the latest version of the ATNF catalog, the pulsar B1055-52 (J1057-5226) is located at a distance of 0.093 kpc. However, the true distance to B1055-52 remains uncertain, varying between 714 and 93 pc depending on the Galactic free-electron density models used: [44] and [43], respectively, despite both using the same dispersion measure. This pulsar is highly energetic, with a spin-down power of $3\times 10^{34}$ erg/s and a spin-down age of 535 kyr. Even if a distance of 714 pc is considered, which would reduce the expected $e^{+}$ flux, the combination of its $\dot{E}$ and $t$ values suggests that B1055-52 would likely still rank among the top 10 most significant sources when compared to the characteristics of the other leading pulsars reported in Table 3. Alongside Geminga and Monogem, B1055-52 is one of the most studied nearby middle-aged pulsars, showing $\gamma$-ray, X-ray, and radio emissions, as highlighted in a recent multi-wavelength study [96]. The pulsar is associated with the point-like $\gamma$-ray source 4FGL J1057.9-5227 in the Fermi-LAT 4FGL catalog [97], showing emissions up to a few GeV. No TeV emissions have been detected thus far. A significant X-ray flux has been observed, including in the recent eRASS:4 catalog obtained by SRG/eROSITA in the 0.2-2.3 keV energy band [98]. Although prominent PWN emission has not been detected in the X-ray band, faint extended emission in a small annulus around the pulsar, possibly linked to its PWN, has been observed by Chandra [99]. The possible detection of a $\gamma$-ray halo around this pulsar is very challenging with GeV data from Fermi-LAT because, due to the proximity of the source, the halo could span several tens of degrees or even cover the entire sky. Detecting a signal in X-ray data is also likely prohibitive due to the typically small field of view of detectors. However, at multi-TeV energies, this source can be detected as an extended source a few degrees in size. It is not within the field of view of HAWC or LHAASO, but it can be detected by HESS.
@@
The potential significance of the "three musketeers" in explaining the $e^{+}$ excess was recently reiterated in [108]. However, their study was limited to a two-zone diffusion scenario and did not account for statistical variation of the injection parameters, nor did it include a fit to the data.